OpenStudy (pphalke):

Malcolm paid an average of $31 per tie for a recent purchase of three ties. The price of one tie was twice as much as another, and the remaining tie cost $24, what were the prices of the other two ties?

OpenStudy (dayakar):

let cost of first tie = x cost of second tie = 2x cost of third tie = 24 average cost = $31 \[\frac{ \sum of the costs of three ties }{ no. of ties } = average\]

OpenStudy (dayakar):

\[\frac{ x+2x+ 24 }{ 3 } = 31\] find x

OpenStudy (mathstudent55):

Another approach: The average price of the three ties was $31. That means all three ties cost 3 * $31 = $93. The first tie cost $24. Subtract the known cost: $93 - $24 = $69 Now you need to find the prices of two ties, knowing that one cost twice the other, and together they cost $69. x + 2x = $69 Solve for x to find the price of the second tie. Then multiply x by 2 to find the price of the third tie.

OpenStudy (pphalke):

@mathstudent55 its 23 because you get 3x=69 then u divide by 3 and get 23 what do I do next

OpenStudy (faiqraees):

24+x+2x = 3 * 31

OpenStudy (pphalke):

what was the price of the less expensive of the two remaining ties? its asking that

OpenStudy (faiqraees):

the value of x is the answer to your current question

OpenStudy (pphalke):

faiqraees I did all that it came out as a decimal that doesn't sound right

OpenStudy (mathstudent55):

Correct. x = 23 is the price of one tie. Now the other tie was 2x, so multiply 23 by 2.

OpenStudy (mathstudent55):

First tie we were told was $24. Second tie, you calculated $23. This was x. Third tie is 2x, so it is simply 2 times $23.

OpenStudy (mathstudent55):

BTW, I did write above what you need to do to find the prices of the other two ties: "x + 2x = $69 Solve for x to find the price of the second tie. Then multiply x by 2 to find the price of the third tie."

OpenStudy (pphalke):

23*2 is 46 @mathstudent55